# Vibration-based damage detection of a trainer jet via multiple input tangential interpolation

### v3

@@ Abstract
Control engineering is a highly developed field, which includes similarly advanced areas like system identification. In structural dynamics, system identification methods are employed for the extraction of modal parameters, such as natural frequencies and mode shapes, from any structure. In turn, these are the main building blocks of vibration-based damage detection. However, traditional comparisons of these parameters are often ambiguous in complex systems, complicating damage detection and assessment. The modified total modal assurance criterion (MTMAC), an index well-known in the field of finite element model updating, is extended to address this challenge and is proposed as a index for damage identification and severity assessment. To support the requirement for precise and robust modal identification of Structural Health Monitoring (SHM), the improved Loewner Framework (iLF), known for its reliability and computational performance, is pioneeringly employed within SHM. Since the MTMAC is proposed solely as a damage identification and severity assessment index, the coordinate modal assurance criterion (COMAC), also a well-established tool, but for damage localisation using mode shapes, is used for completeness. The iLF SHM capabilities are validated through comparisons with traditional methods, including least-squares complex exponential (LSCE) and stochastic subspace identification with canonical variate analysis (SSI-CVA) on a numerical case study of a cantilever beam. Furthermore, the MTMAC is validated against the traditional vibration-based approach, which involves directly comparing natural frequencies and mode shapes. Finally, an experimental dataset from a BAE Systems Hawk T1A trainer jet ground vibration tests is used to demonstrate the iLF and MTMAC capabilities on a real-life, real-size SHM problem, showing their effectiveness in detecting and assessing damage.
@@
Loewner Framework , System Identification , Multi-input multi-output , Experimental Modal Analysis , Structural Health Monitoring , Aeronautical Systems , MTMAC , Tangential Interpolation
@@
{highlights}
A system identification-driven structural health monitoring (SHM) method is proposed;
The two-step method involves modal parameter extraction and evaluation via a combined metric;
The improved Loewner Framework is used to extract the modal parameters, a first for SHM;
The modified total modal assurance criterion is used, as a first, as a standalone SHM index;
The method is validated on a numerical system and applied to a full-size real-life trainer jet.
@@ 1 Introduction
In the general control engineering domain, the field of system identification is critical as a way to extract information and/or obtain models from simulation, experimental, or real data [1]. For example, in structural dynamics, system identification methods are used for the extraction of modal parameters [2], such as natural frequencies ($\omega_{n}$), damping ratios ($\zeta_{n}$), and mode shapes ($\bm{\phi}_{n}$), from linear time invariant systems (LTIs) [3]. Modal parameters have always been the backbone of vibration-based Structural Health Monitoring (SHM) [4]. Hence, there is a continuous search for novel and efficient methods for extracting modal parameters, such as those presented in [5], alongside more refined indices [6] and analysis methods [7]. The former are necessary to move away from direct comparison of modal parameters -– a notoriously challenging task for higher-dimensional systems. For this reason, indices such as the modal assurance criterion (MAC) [8], which helps condense information to make it usable for SHM, have achieved considerable success over the years. On the other hand, the latter is not restricted to direct methods and can include indirect methods, such as model updating-based approaches [9, 10], as well as data-driven techniques [11]. This continuous stream of emerging methods and strategies is also fuelled by the fact that most engineering fields require some sort of damage monitoring or identification strategy. Some of the more well-known applications are infrastructures, such as roads [12] and wind turbines [13], and systems, such as fuel pumps, [14].
@@
Damage can be defined as any structural change in a system that compromises or affects its operational capacity [15]]. The process of implementing a damage detection strategy for aerospace [16], civil [11], or mechanical engineering [17] systems is commonly referred to as SHM. In the case of vibration-based SHM, modal parameters, such as $\omega_{n}$, $\zeta_{n}$, and $\bm{\phi}_{n}$, are frequently selected as damage indicators due to their ease of identification from vibrational data [18] and their direct relationship with the mass and stiffness of the structure. These parameters can serve various purposes, such as assessing damage severity and/or localisation. $\omega_{n}$ are particularly effective for evaluating damage severity, but less reliable for localisation since small frequency changes are often masked by environmental and operational variations [19]. Nevertheless, they were successfully applied in [20] for the SHM of a masonry bridge structure using single input and multi output (SIMO) data. Conversely, $\bm{\phi}_{n}$ are better suited for damage localisation, typically achieved by detecting deviations between baseline and damaged values. Still, they can also be helpful for damage severity assessment. For example, [21] proposed a two-step procedure for damage detection on beam-like structures using $\bm{\phi}_{n}$. $\zeta_{n}$ are hardly ever used as a standalone damage indicator due to their strong dependence on non-structural factors, which makes it difficult to distinguish changes in damping caused by damage from those due to other influences [22]. Recent advances in vibration-based SHM have focused on civil infrastructure, particularly bridges [23] and wind turbines [24], with a growing interest in addressing challenges such as uncertainty quantification, noise robustness, and operational variability [25]. Advanced methodologies, such as the Cepstrum-Informed Attention-Based Network, have been introduced to model structural dynamics efficiently while maintaining physical interpretability, using compact damage-sensitive features derived from cepstral coefficients [26]. Bayesian modal identification approaches have been developed to incorporate ambient and seismic response data, enabling robust modal parameter estimation with reduced computational complexity [27]. Additionally, digital twin frameworks have established themselves as effective tools for predictive maintenance, offering unambiguous correlations between dynamic behaviours and structural conditions, such as under-foundation scour in bridges [28]. For wind turbine rotors, operational modal analysis techniques have been adapted to identify localised stiffness reductions and anisotropic effects using time-periodic approximations and Gaussian residual tests [29]. Furthermore, machine learning methods, including convolutional neural networks, have been successfully applied to classify damage scenarios in lightweight bridge structures, addressing the complexities of coupled dynamic responses from moving loads [30]. Lastly, population-based SHM has demonstrated the potential to transfer structural behaviour knowledge from homogeneous building populations to individual targets, offering a scalable solution for damage detection across nominally identical structures [31].
@@
Nevertheless, an unequivocal solution for SHM does not yet exist, as most methods and strategies are either underdeveloped or not fully codified in national or international standards [32]. In addition, developing new methods – or refining existing ones – is a challenge due to the lack of diverse datasets from the system experimental setup perspective. In fact, most case studies and benchmarks for vibration-based SHM focus on single-input multi-output (SIMO) [33] or output-only data [34] rather than multiple input and multiple output (MIMO). This is critical as MIMO testing can offer further insight into modes not fully identified from SIMO or OMA campaigns, such as out-of-plane dominant motion in flexible wings [3] or civil infrastructure. A well-known example is the Z24 bridge case study, where the weakly excited lateral modes are the most difficult to identify [35]. Despite this difficulty, [36] showed that the first lateral mode was the most indicative of damage occurrence, which can be extrapolated as an argument to use identification of small amplitude vibration modes to enhance the precision of vibration-based SHM. While some efforts were made at the beginning of the 2000s for MIMO-based bridge monitoring [37, 38], these were not followed by further advances or testing campaigns. Additionally, there is a general lack of aeronautically relevant examples for SHM, with the notable exception of [39, 40]. However, this issue was recently addressed by the Laboratory for Verification and Validation (LVV) at the University of Sheffield, which conducted a thorough vibration testing campaign for a full-scale trainer jet aircraft [41] and has released the dataset [42]. This includes vibration data from different input signals, including simulated and real damaged cases in a MIMO configuration, offering a thorough standard for future identification and damage detection methods. Addressing the fact that damage assessment and localisation through modal parameters direct comparison for complex structures is difficult at best. Thus, a two-step modal & damage identification framework is proposed. The first step uses the improved Loewner Framework (iLF) [43] to identify the modal parameters in the system, while the second step uses the modified total modal assurance criterion (MTMAC) [44] as a damage assessment index based on the identified modal parameters. The iLF, a recently developed computationally efficient MIMO modal parameter identification technique, is applied in this article to the trainer jet aircraft dataset mentioned above and, beforehand, validated on a numerical case study. The work reported here represents the first instance in the literature where iLF is applied for modal parameter-based SHM purposes. To highlight the capabilities of the iLF, the results are compared to those from other industrially relevant methods, such as least-squares complex exponential (LSCE) [45] and stochastic subspace identification with canonical variate analysis (SSI-CVA) [2]. The former is a frequency domain method (as is the iLF), while the latter is an output-only technique in the time domain. In addition, the MTMAC (modified total modal assurance criterion) is shown to be a valid index for damage assessment. It is highlighted that this criterion was used before, together with MACFLEX (modal flexibility assurance criterion) [46], but is here proposed as a standalone damage assessment index. Further, the well-known Coordinate Modal Assurance Criterion (COMAC) [47] is used for damage localisation to ensure the completeness of the damage assessment procedure. The rationale is that the COMAC identifies which degrees-of-freedom (DoFs) measurements contribute to a low value of MAC [48]; hence, a low value of COMAC corresponds to a local $\bm{\phi}_{n}$ trajectory deviation, which is a typical sign of damage. This is supported, among many other examples, by the relevant literature on damage localisation on composite laminates [49], reinforced concrete beams [50], and retrofitted unreinforced masonry walls [51].
@@
    The framework (iLF and MTMAC) is applied to a comprehensive dataset obtained from a full-scale trainer jet aircraft.
@@ 2 Methodology
The following subsections outline the iLF theoretical background (Section 2.1) and the SHM tools considered (Section 2.2). Then, in Section 2.3, the proposed approach of using the iLF as a valid tool for SHM purposed is discussed
@@ 2.1 The Loewner Framework
Before moving forward, a brief technical presentation of the LF, on which the iLF computational implementation is based, is given, starting with the description of the Loewner matrix $\mathbb{L}$. For a row array of pairs of complex numbers ($\mu_{j}$,${v}_{j}$), $j=1$,…,$q$, and a column array of pairs of complex numbers ($\lambda_{i}$,${w}_{j}$), $i=1$,…,$\rho$, such that $\lambda_{i}$, $\mu_{j}$ distinct, the associated $\bm{\mathbb{L}}$ can be defined:
@@
The indices $\lambda_{i}$ and $\mu_{j}$ represent, respectively, the values at which the right ($\mathbf{R}$) and left ($\mathbf{L}$) subsets are sampled to produce the approximation $\mathbf{H}(s)$, while $\rho$ and $q$ correspond to their dimension. The vectors and denote the right and left tangential general directions, typically chosen randomly in practice [57], while the values and represent the right and left tangential data.
The rational interpolation problem is resolved when the transfer function $\mathbf{H}(s)$, associated with realisation $\mathbf{\Sigma}$ in Equation 2, is linked to $\mathbf{w}_{i}$ and $\mathbf{v}_{i}$ in such a way that the Loewner pencil satisfies Equation 6:
$\begin{split}\mathbf{H}(\lambda_{i})\mathbf{r}_{i}=\mathbf{w}_{i},\>j=1.\dots,\rho\\ \mathbf{l}_{i}\mathbf{H}(\mu_{j})=\mathbf{v}_{j},\>i=1,\dots,q\end{split}$ (6)
@@
Now, given a set of points $Z=\{z_{1},\dots,z_{N}\}$ in the complex plane, the corresponding value of $\mathbf{H}(s)$ can be separated into left $\lambda_{i=1...\rho}$ and right $\mu_{i=1...q}$ data with $N=q+\rho$:
@@
This leads to the following system matrix assignment::
@@
and likewise, $-\mathbf{Y}\mathbf{A}+\mathbf{M}\mathbf{Y}\mathbf{E}=\mathbf{L}\mathbf{C}$ such that $\mathbf{X}$ and $\mathbf{Y}$ are the generalised controllability and observability matrices for the system $\mathbf{\Sigma}$ with $\mathbf{D}=0$. After verifying the right and left interpolation conditions, the Loewner realisation for redundant data is as follows:
@@
2.2 Damage sssessment, quantification, and localisation
As hinted in Section 1, damage can be assessed by identifying a difference in $\omega_{n}$ between the baseline and a successive structural state. Furthermore, quantifying this difference can pinpoint to the severity level of the damage, e.g., higher differences might be related to a more severe damage state. On the other hand, localisation is achieved by comparing the baseline and damaged $\bm{\phi}_{n}$, where deviations from the baseline shape indicate the damage location. These methodologies are well-established in the literature. Nonetheless, these deviations can also be used for assessment and quantification since a more significant deviation corresponds to a more severe damage scenario. Nevertheless, these comparisons become challenging to visualise and interpret for complex structures. Hence, in this work, the complement of the well-known MTMAC [44] is proposed for damage assessment, to pair the $\omega_{n}$ and $\bm{\phi}_{n}$ contributions to damage assessment and quantification:

### v4

# Gabriele Dessena Marco Civera Andrés Marcos Bernardino Chiaia Oscar E. Bonilla-Manrique

@@ Abstract
The problem of damage detection and identification is of interest for many aerospace and aeronautical engineering systems. However, relevant literature mostly focuses on subsystems and parts, rather than full airframes. In structural dynamics, modal parameters, such as natural frequencies and mode shapes, from any structure are the main building blocks of vibration-based damage detection. However, traditional comparisons of these parameters are often ambiguous in complex systems, complicating damage detection and assessment. The modified total modal assurance criterion (MTMAC), an index well-known in the field of finite element model updating, is extended to address this challenge and is proposed as an index for damage identification and severity assessment. To support the requirement for precise and robust modal identification of Structural Health Monitoring (SHM), the improved Loewner Framework (iLF), known for its reliability and computational performance, is pioneeringly employed within SHM. Since the MTMAC is proposed solely as a damage identification and severity assessment index, the coordinate modal assurance criterion (COMAC), also a well-established tool, but for damage localisation using mode shapes, is used for completeness. The iLF SHM capabilities are validated through comparisons with traditional methods, including least-squares complex exponential (LSCE) and stochastic subspace identification with canonical variate analysis (SSI-CVA) on a numerical case study of a cantilever beam. Furthermore, the MTMAC is validated against the traditional vibration-based approach, which involves directly comparing natural frequencies and mode shapes. Finally, an experimental dataset from a BAE Systems Hawk T1A jet trainer ground vibration test is used to demonstrate the iLF and MTMAC capabilities on a real-life, real-size SHM problem, showing their effectiveness in detecting and assessing damage.
@@
Loewner Framework , Jet Trainer , Multi-input multi-output , Damage Detection , Structural Health Monitoring , Aeronautical Structure , MTMAC , Tangential Interpolation
journal: Aerospace Science & Technology
@@ 1 Introduction
Damage can be defined as any structural change in a system that compromises or affects its operational capacity [1]. The process of implementing a damage detection strategy for aerospace [2], civil [3], or mechanical engineering [4] systems is commonly referred to as Structural Health Monitoring (SHM).
@@
In the case of vibration-based SHM, modal parameters, such as natural frequencies ($\omega_{n}$), damping ratios ($\zeta_{n}$), and mode shapes ($\bm{\phi}_{n}$), are frequently selected as damage indicators due to their ease of identification from vibrational data [5] and their direct relationship with the mass and stiffness of the structure. These parameters can serve various purposes, such as assessing damage severity and/or localisation. $\omega_{n}$ are particularly effective for evaluating damage severity, but less reliable for localisation since small frequency changes are often masked by environmental and operational variations [6]. Nevertheless, they were successfully applied in [7] for the SHM of a masonry bridge structure using single-input and multiple-output (SIMO) data. Conversely, $\bm{\phi}_{n}$ are better suited for damage localisation, typically achieved by detecting deviations between baseline and damaged values. Still, they can also be helpful for damage severity assessment. For example, [8] proposed a two-step procedure for damage detection on beam-like structures using $\bm{\phi}_{n}$. $\zeta_{n}$ are hardly ever used as a standalone damage indicator due to their strong dependence on non-structural factors, which makes it difficult to distinguish changes in damping caused by damage from those due to other influences [9].
Beyond these traditional strategies, more advanced approaches have emerged. Recent advances in vibration-based SHM have focused on civil infrastructure, particularly bridges [10] and wind turbines [11], with a growing interest in addressing challenges such as uncertainty quantification, noise robustness, and operational variability [12]. Furthermore, clustering algorithms, such as the Density-based Spatial Clustering of Applications with Noise (DBSCAN), have gained attention for identifying aeronautically relevant structures. Notable applications are the output-only modal identification of a helicopter blade in [13] and an aircraft model in [14]. Advanced methodologies, such as the Cepstrum-Informed Attention-Based Network, have been introduced to model structural dynamics efficiently while maintaining physical interpretability, using compact damage-sensitive features derived from cepstral coefficients [15]. Bayesian modal identification approaches have been developed to incorporate ambient and seismic response data, enabling robust modal parameter estimation with reduced computational complexity [16]. Additionally, digital twin frameworks have established themselves as effective tools for predictive maintenance, offering unambiguous correlations between dynamic behaviours and structural conditions, such as under-foundation scour in bridges [17]. For wind turbine rotors, operational modal analysis techniques have been adapted to identify localised stiffness reductions and anisotropic effects using time-periodic approximations and Gaussian residual tests [18]. Furthermore, machine learning methods, including convolutional neural networks, have been successfully applied to classify damage scenarios in lightweight bridge structures, addressing the complexities of coupled dynamic responses from moving loads [19]. Moreover, as shown in [20], traditional guided wave SHM can be enhanced for operational damage detection in aeronautical structures using Gaussian mixture models and their probability function for damage detection and localisation. Lastly, population-based SHM has demonstrated the potential to transfer structural behaviour knowledge from homogeneous building populations to individual targets, offering a scalable solution for damage detection across nominally identical structures [21].
Nevertheless, an unequivocal solution for SHM does not yet exist, as most methods and strategies are either underdeveloped or not fully codified in national or international standards [22]. Furthermore, in aerospace and aeronautical engineering, most applications focus on a deep analysis of individual subsystems, like fuel pumps [23] and tanks [24], or parts, like blades [25], attachment lugs [26], aircraft engine brackets [27, 28, 29], and spars [30], rather than full or partial airframes, despite very early efforts [31].
In addition, developing new methods – or refining existing ones - is a challenge due to the lack of diverse datasets from the system experimental setup perspective. In fact, most case studies and benchmarks for vibration-based SHM focus on SIMO [32] or output-only data [33] rather than multiple input and multiple output (MIMO). This is critical as MIMO testing can offer further insight into modes not fully identified from SIMO or OMA campaigns, such as out-of-plane dominant motion in flexible wings [34] or civil infrastructure. A well-known example is the Z24 bridge case study, where the weakly excited lateral modes are the most difficult to identify [35]. Despite this difficulty, [36] showed that the first lateral mode was the most indicative of damage occurrence, which can be extrapolated as an argument to use the identification of small amplitude vibration modes to enhance the precision of vibration-based SHM.
While some efforts were made at the beginning of the 2000s for MIMO-based bridge monitoring [37, 38], these were not followed by further advances or testing campaigns. Additionally, there is a general lack of aeronautically relevant examples for SHM, with the notable exception of [39, 40]. However, this issue was recently addressed by the Laboratory for Verification and Validation (LVV) at the University of Sheffield, which conducted a thorough vibration testing campaign for a full-scale jet trainer aircraft [41] and has released the dataset [42]. This includes vibration data from different input signals, including simulated and real damaged cases in a MIMO configuration, offering a thorough standard for future identification and damage detection methods.
Finally, coming back to the basics of vibration-based SHM, there is a third, and fundamental, critical point: the reliability of the identified modal parameters.
In this regard, system identification (SI) offers critical approaches to extract information and/or obtain models from simulation, experimental, or real data [43]. For example, in structural dynamics, SI methods are used for the extraction of modal parameters [44], such as natural frequencies ($\omega_{n}$), damping ratios ($\zeta_{n}$), and mode shapes ($\bm{\phi}_{n}$), from linear time invariant systems (LTIs) [34]. Modal parameters have always been the backbone of vibration-based SHM [45]. Hence, there is a continuous search for novel and efficient methods for extracting modal parameters, such as those presented in [46], alongside more refined indices [47] and analysis methods [48]. The former are necessary to move away from direct comparison of modal parameters – a notoriously challenging task for higher-dimensional systems. For this reason, indices such as the modal assurance criterion (MAC) [49], which helps condense information to make it usable for SHM, have achieved considerable success over the years. On the other hand, the latter is not restricted to direct methods and can include indirect methods, such as model updating-based approaches [50, 51], as well as data-driven techniques [3].
@@
Given all these challenges - lack of definitive damage indices, MIMO experimental validation, and efficient SI algorithms - We propose a two-step modal & damage identification framework. The first step uses the improved Loewner Framework (iLF) [52] to identify the modal parameters from a MIMO system, while the second step uses the modified total modal assurance criterion (MTMAC) [53] as a damage assessment index based on the identified modal parameters. The iLF, a recently developed computationally efficient MIMO modal parameter identification technique, is applied in this article to the jet trainer aircraft dataset mentioned above and, beforehand, validated on a numerical case study. The work reported here represents the first instance in the literature where iLF is applied for modal parameter-based SHM purposes. To highlight the capabilities of the iLF, the results are compared to those from other industrially relevant methods, such as least-squares complex exponential (LSCE) [54] and stochastic subspace identification with canonical variate analysis (SSI-CVA) [44]. The former is a frequency domain method (as is the iLF), while the latter is an output-only technique in the time domain. In addition, the MTMAC (modified total modal assurance criterion) is shown to be a valid index for damage assessment. It is highlighted that this criterion was used before, together with MACFLEX (modal flexibility assurance criterion) [55], but is here proposed as a standalone damage assessment index. Further, the well-known Coordinate Modal Assurance Criterion (COMAC) [56] is used for damage localisation to ensure the completeness of the damage assessment procedure. The rationale is that the COMAC identifies which degrees-of-freedom (DoFs) measurements contribute to a low value of MAC [57]; hence, a low value of COMAC corresponds to a local $\bm{\phi}_{n}$ trajectory deviation, which is a typical sign of damage. This is supported, among many other examples, by the relevant literature on damage localisation on composite laminates [58], reinforced concrete beams [59], and retrofitted unreinforced masonry walls [60].
Consequently, the main aim of this study is the introduction of a condensed index for damage detection using the MTMAC, which leads to the following objectives:
  1. 1.
    Validate the iLF precision and accuracy for modal parameter-based damage detection via direct comparison on a numerical model;
  2. 2.
    Show that the MTMAC can be a valid and accurate damage index on a numerical model;
  3. 3.
    Validate the proposed combined iLF and MTMAC approach on the numerical data;
  4. 4.
    Apply the damage detection strategy on a complex aeronautical structure: A BAE Systems Hawk T1A airframe
Being able to use the proposed combined approach is extremely important for complex systems, such as a full trainer jet, because modal parameters are difficult to monitor directly, and a merged index, like the proposed MTMAC, makes differentiating between healthy and damaged cases a more straightforward task. Moreover, it is noted that the proposed method is formulated for linear time-invariant dynamics, as is customary in modal-based SHM. This is consistent with the ground vibration test dataset, which was acquired under low-amplitude excitations ensuring quasi-linear structural behaviour. Nevertheless, although outside the scope of this work, the interested reader is referred to [61] for the theoretical background that allows the nonlinear modal analysis of a full aircraft, a Morane-Saulnier Paris, shown in [62].
@@
    The framework (iLF and MTMAC) is applied to a comprehensive dataset obtained from a full-scale jet trainer aircraft.
@@ 2 Methodology
The following subsections outline the iLF theoretical background (Section 2.1) and the SHM tools considered (Section 2.2). Then, in Section 2.3, the proposed approach of using the iLF as a valid tool for SHM purposes is discussed.
@@ 2.1 The Loewner Framework
[…]
@@
The indices $\lambda_{i}$ and $\mu_{j}$ represent, respectively, the values at which the right ($\mathbf{R}$) and left ($\mathbf{L}$) subsets are sampled to produce the approximation $\mathbf{H}(s)$, while $\rho$ and $q$ correspond to their dimension. The vectors $\mathbf{w}_{i}$ and $\mathbf{v}_{j}$ denote the right and left tangential general directions, typically chosen randomly in practice [67], while the values $\mathbf{r}_{i}$ and $\mathbf{l}_{j}$ represent the right and left tangential data.
The rational interpolation problem is resolved when the transfer function $\mathbf{H}(s)$, associated with realisation $\mathbf{\Sigma}$ in Equation 2, is linked to $\mathbf{w}_{i}$ and $\mathbf{v}_{j}$ in such a way that the Loewner pencil satisfies Equation 6:
$\begin{split}\mathbf{H}(\lambda_{i})\mathbf{r}_{i}=\mathbf{w}_{i},\>i=1.\dots,\rho\\ \mathbf{l}_{j}\mathbf{H}(\mu_{j})=\mathbf{v}_{j},\>j=1,\dots,q\end{split}$ (6)
@@
Now, given a set of points $Z=\{z_{1},\dots,z_{N}\}$ in the complex plane, the corresponding value of $\mathbf{H}(s)$ can be separated into left $\lambda_{i=1...\rho}$ and right $\mu_{j=1...q}$ data with $N=q+\rho$:
@@
This leads to the following system matrix assignment:
@@
and likewise, $\mathbf{Y}\mathbf{A}+\mathbf{M}\mathbf{Y}\mathbf{E}=\mathbf{L}\mathbf{C}$ such that $\mathbf{X}$ and $\mathbf{Y}$ are the generalised controllability and observability matrices for the system $\mathbf{\Sigma}$ with $\mathbf{D}=0$. After verifying the right and left interpolation conditions, the Loewner realisation for redundant data is as follows:
@@
2.2 Damage assessment, quantification, and localisation
As hinted in Section 1, damage can be assessed by identifying a difference in $\omega_{n}$ between the baseline and a successive structural state. Furthermore, quantifying this difference can pinpoint the severity level of the damage, e.g., higher differences might be related to a more severe damage state. On the other hand, localisation is achieved by comparing the baseline and damaged $\bm{\phi}_{n}$, where deviations from the baseline shape indicate the damage location. These methodologies are well-established in the literature. Nonetheless, these deviations can also be used for assessment and quantification since a more significant deviation corresponds to a more severe damage scenario. Nevertheless, these comparisons become challenging to visualise and interpret for complex structures. Hence, in this work, the complement of the well-known MTMAC [53] is proposed for damage assessment, to pair the $\omega_{n}$ and $\bm{\phi}_{n}$ contributions to damage assessment and quantification: